An angle @ is drawn in standard position. The point (5,-12) is located on the terminal side of @. Find the exact values of the six trigonometric functions of @.
Answer by Alan3354(69443)

An angle @ is drawn in standard position. The point (5,-12) is located on the terminal side of @. Find the exact values of the six trigonometric functions of @.
---------------
r = sqrt(x^2 + y^2) = 13
===========
sin = y/r
cos = x/r
tan = y/x
---
cot = 1/tan
sec = 1/cos
csc = 1/sin
===============
The point is in Q4, so adjust the signs (plus or minus) as needed.